Home‎ > ‎Reports‎ > ‎Crew102‎ > ‎

day10


Crew 102 Log Book for April 4, 2011

Photos of the day





Commander's Check-In Report

Franco Carbognani Reporting

Crew Physical Status:  John went today to the Medical Clinic and he was diagnosed an eye infection plus throat and ears inflammation. He has been prescribed the appropriate medicaments. No particular concerns. Rest of the crew is doing fine.

Time Departed/Returned from EVA: 14:30 - 15:35, 16:45 - 20:05

Brief Narrative of Field Mission Results:

Repeated part of the soil permeability experiment to obtain more accurate data

Mapped 44 points in the border between bio-erosion in Griphea Arcuata and corrosion.


EVA Data/Interpretations:

Cyanobacteria samples analysis ongoing

We have been able to find a way to download data from the GPS device available onsite (This may be save us some network bandwidth used to download GPS data from the  smart-phones)

Engineering/Hab Maintenance:

Main generator running fine after provisional fixes

Downstairs heating is not working. The furnace serving is stopped and is reporting an error. With the use of the wall mounted heater temperatures remain in good range.

Plans for Tomorrow:

At 09:00 AM the filming crew from Arte Television will be with us again.

A "filming friendly" ATV EVA in the most suitable cyanobacteria site near by.

Another EVA dedicated to Griphea Arcuata bioerosion.

Short EVA around the Hab dedicated to corrosion inspection and sample collection for optical microscopy analysis.

Proceeding on bringing back gray water system pipeline.

Continue on ongoing psychological and physiological studies.

Support Requested:

Don Lusco for:

1) Final fixes on power generator.

2) Fix downstairs furnace

3) Refill water trailer tank


Commander's Report

Franco Carbognani Reporting

At 10:30 AM the filming crew from Arte Television arrived on-site. We have shortly discussed with them the details of the documentary they plan to film. We have started with a general overview of the MDRS and then the 2 EVAs were filmed plus the interview of one crew member.

The furnace serving downstairs is stopped and is reporting an error.

Using as a backup the available wall mounted heater.

Radiotelescope: Still hoping the radiotelescope calibrator could arrive on-site within our rotation.


Engineering Report

Franco Carbognani Reporting

Generator/Electricity:

Kitty (Diesel Generator) (Used? Yes/No): Yes

Honey (Gas Generator) (Used? Yes/No): No

Battery Charge Level:

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Kitty provisional repair seems holding. The cooling fan is running smoothly as the whole Kitty engine.


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 0.62

Propane Tank (Full/0.75/0.5/0.25/Empty): 0.5

Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.6

Notes/Comments:


Oil:

Kitty Oil Quantity (# of Quarts in storage): 4/4

Honey Oil Quantity (# of Quarts in storage): 0/4, to be refilled

ATV Oil Quantity (# of Quarts in storage): all 4/4

Car Oil Quantity (# of Quarts in storage): 3/4


Water Status:

Outside Potable Water Tank Level (inches from bottom): 23

Trailer Potable Water Tank Level (inches from bottom): 0

Hab Potable Water Tank Level (inches from bottom): 35

Potable Water Meter Reading: 44545,0

In to GreenHab Meter Reading:

Notes/Comments:

The Trailer Potable Water Tank is now empty and ready for Don to be refilled.


GreenHab:

Condition of plants in Tank 1:

Condition of plants in Tank 2:

Condition of plants in Tank 3:

Greenhab Notes/Comments:

Gray Water System: The Trickle Reservoirs are now filled with gray water. A tablespoon of proline bacteria has been pured on both of them and the pumps restarted.


Observatory:

Telescope Used? (Yes/No):No

Observatory Notes/Comments:

Boxes containing the telescope mount and controller are ready into the Observatory in order to be handed over to Don


Transportation:

Hab Car (NOT RENTAL)

Used (yes/no): yes

Oil Checked (yes/no):yes

Tire Status: good

Odometer at end of day: 123903.8

Notes/Comments on Hab Car:


ATV 1 (Opportunity):

Used: yes

Oil Checked: n

Fuel Consumed: 1/2 liter

Tire Status: inflated

Comments:


ATV 2 (Spirit):

Used: yes

Oil Checked: no

Fuel Consumed: 1/2 liter

Tire Status: inflated

Comments:


ATV 3 (Viking I):

Used: no

Oil Checked: n

Fuel Consumed: none

Tire Status: good

Comments:


ATV 4 (Viking II):

Used: no

Oil Checked: n

Fuel Consumed: none

Tire Status: good

Comments: this ATV does not start (this one and not Spirit as previusly wrongly reported)


Heating and Ventilation:

Thermostat setting upstairs (Farenheit): 65

Actual temperature upstairs (Farenheit): 72

Thermostat setting downstairs (Farenheit):not relevant

Actual temperature downstairs (Farenheit): 65

Heating and Ventilation Notes/Comments:

Downstairs furnace is not working and reporting an error. Don acknowledged the problem. He thinks the problem is the furnace fan not working properly. We are using as a backup the wall mounted heater which is working fine. With current temperatures the situation is manageable.


Computers/Networking Infrastructure:

Notes/comments on internet/computers?:


General Engineering Notes/Comments:


Science Report (Human factors/Psychology)

Dr. John E. Deaton Reporting

Psychology tests (Stress, Leadership Evaluation, Exploration Growth Scale, and Emotional Scale) are progressing according to original plan and protocol.    Human performance tests are likewise progressing according to schedule and consist of multiple tests to measure sleepiness ratings, human performance (including workload measurements obtained after each EVA) and cognitive processing.  These tests will be repeated every other day until the completion of the mission, April 8.



Health and Safety Officer's Report

Dr. John E. Deaton Reporting

Health and Safety Officer and another crew member  temporarily broke sim to travel to the nearest medical clinic in Bicknell, UT as to undergo tests for conjunctivitis (eye infection).  Dr. Czarnik diagnosed the condition after submitting a photograph of both eyes, and recommended antibacterial eye drops. Crew returned from Bicknell clinic and indeed, Dr. Czarnik was correct.  Personnel at the clinic diagnosed an eye infection, nose, throat, and ear inflammation, and prescribed cough medicine to control congestion and cough, and an antibiotic to reduce swelling in the ENT.  It is the opinion of the HSO that the medical teleconferencing system at the MDRS works extremely well given the most rapid response from the Flight Surgeon in attendance, and the system has certainly avoided what could have been more severe consequences.


Other crew members have not reported anything out of the ordinary, with the exception of some mild intestinal discomfort.


EVA 10 Report

Kavya k.Manyapu Reporting

Time of Departure:  2.30 pm

Time of Arrival: 3.35 pm

Crew Members: K.Manyapu, F.Carbognani

Site Location:

Latitude:     38ᵒ24’23”

Longitude:  -110ᵒ47’30”

Altitude: 4500 feet

Total Distance: 0.4 miles

Transit Mode: By foot

Objectives: The purpose of this EVA was to perform soil sampling and to conduct experiments to measure the soil permeability and soil porosity for the various types of soils in this area. This experiment is a supplementary to the EVA that was performed last week for the same purpose in order to collect more data to validate the results from the experiments.  A science report will follow once the data has been analyzed.

Results:  The crew has been successful in obtaining data that was similar to the first experiment performed a few days ago.

Lessons Learned:

A humidity and temperature sensor for subsurface geology experiments will be very helpful to have in the Hab for future crew and experiments.


EVA 11 Reporting

Yuval Brodsky Reporting

Time: 1645 - 2005

Crew Members: Y. Brodsky (EVA CDR), A. Kapoglou

Site Location: Northwestern portion of Radio Ridge

Transit Mode: ATV

Objectives: Locating bio-erosion on Gryphea arcuata and corrosion, and mapping both phenomena (and the boundary between both) over 43 pre-defined waypoints.

Results: Mapping was successfully completed over all pre-defined waypoints. Boundary between both phenomena was apparent.

Lessons Learned: N/A.


Journalist Report

Kavya Manyapu Reporting

Voila, the French were here today! Crew 102 is being visited by a French TV Channel called ARTE for the next 3 days. The crew has been conducting its regular activities while the French crew has been filming them. The crew is very excited to show the world what goes on at MDRS and how motivated Mars Society, its volunteers and crew are in promoting Human Space Exploration.  After a brief introduction by the commander with all aspects of the MDRS Simulation to the French crew, the crew was back again busy conducting the experiments of the day.

The first EVA of the day was commanded by K.Manyapu and was accompanied by F.Carbognani. A.Kapoglou served as the cap-com. The EVA was a supplementary to the soil sampling for the fuel cell project that was conducted last week by K.Manyapu in order to obtain more data and validate the previous EVA results. The second EVA commanded by bio-erosion and corroded rock samples. K.Manyapu served as the cap-com for the mission.

Y.Brodsky and accompanied by A.Kapoglou was a mission to find

One crew was treated for his infection and is now in better health. For the rest of the day the crew has been busy analyzing data.

The crew has another 5 days left to analyze, finalize and wrap up the experiments while conducting some outreach activity via the French channel.

Keep watching crew 102 in the live webcams to get a closer look of their life on Mars.


Chef's Report

Lara Vimercati Reporting

1a) Today was a non-cooking day

1b) Was today a special day celebrated at the main meal with special

food or activity? If so, what was special about today and what food(s)

were prepared to mark the celebration?

No special celebration today

2) List the foods served at today's main meal (usually the evening meal), giving full names of each.

Forever young macaroni and cheese

Navy Beans soup

Boiled vegetables

Dry fruit

3) List any main meal foods not finished by the crew (leftovers stored for later use or discarded)

Some of the boiled vegetables

4) Recipes for anything more complicated than following package directions exactly, or rehydrating.

N/A

5) Lessons learned (if any) N/A

6) Comments/questions for Kim and Jean, the food study investigators

The crew did not receive this meal very well. Discontent was expressed by the crew because of low variety in the non-cooking items.